OpenStudy (anonymous):

ft 10 ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on a bridge. The horizontal distance that spans is 2ft longer than the height that reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace.

OpenStudy (anonymous):

solve \[x^2+(x+2)^2=10^2\]

OpenStudy (anonymous):

\[x^2+x^2+4x+=100\] \[2x^2+4x=100\] \[x^2+2x=50\] \[(x+1)^2=51\] \[x+1=\sqrt{51}\] \[x=-1+\sqrt{51}\]

OpenStudy (gw2011):

I get an answer of: vertical=6 ft and horizontal=8 ft. Here is how I got the answers: Let x=Vertical Distance Let x+2=Horizontal Distance a^2+b^2=c^2 (x^2)+(x+2)^2=(10)^2 x^2+x^2+4x+4=100 2x^2+4x+4=100 2x^2+4x-96=0 (2x+16)(x-6)=0 2x+16=0 2x=-16 x=-8 This is not a solution because you can't have a negative number x-6=0 x=6 This is the vertical distance x+2=6+2=8 This is the horizontal distance